St. Benedict Hospital has three primary revenue producing departments (Inpatient, Outpatient, Clinic) with the following revenue and cost projections. In order to better determine the departments’ overall true cost and profit margin, management wants to allocate service department costs as overhead allocations to these departments.               
PROJECTED REVENUES AND COSTS PER DEPARTMENT                  
                  
Revenues                  
Inpatient Services   $18,500,000               
Outpatient Services   32,000,000               
Clinic Services   12,250,000               
Total revenues   $62,750,000               
                  
Direct Costs                  
Inpatient Services   $8,250,000               
Outpatient Services   16,500,000               
Clinic Services   8,500,000               
Total costs   $33,250,000               
                  
Service Department Costs                  
Financial Services   $4,000,000               
Facilities   7,500,000               
Housekeeping   5,600,000               
Administration   8,500,000               
Total overhead costs   $25,600,000               
Total Costs   $58,850,000               
Projected Profit   $3,900,000               
                  
Management considered various cost drivers and made the determiniation to use the following as the most relevant for each service department:                  
                  
                  
Department   Cost Driver              
Financial Services   Patient revenue              
Facilities   Square feet              
Housekeeping   Housekeeping hrs              
Administration   Salary dollars              
                  
Utlization information for allocations:                  
Cost Driver   Inpatient   Outpatient   Clinic   Total Utilization  
Square Feet   210,000    180,000    65,000    455,000   
Housekeeping Hours   80,000    58,000    26,000    164,000   
Salary Dollars   $6,500,000    $12,500,000    $5,500,000    $24,500,000   
                  
                  
Based upon this information, complete the following, filling in each highlighted cell. Be sure to use formulas to show your work.                  
                  
                  
Use Exhibit 6.6 as a guideto complete this table and calculate the allocation rate.                  
Department   Cost Pool Amount   Cost Driver   Total Utilization   Allocation Rate  
Financial Services                  
Facilities                  
Housekeeping                  
Administration                  
                  
Use Exhibit 6.7 as a guide to complete the indirect cost allocations to each patient service area.                  
Indirect Cost Allocations   Inpatient   Outpatient   Clinic   Total  
Financial Services                  
Facilities                  
Housekeeping                  
General Administration                  
Total Indirect Costs                  
                  
Using Exhibit 6.8 as a guide, complete the P&L projections below.                  
P&L                  
Revenue                  
Direct Costs                  
Indirect Costs                  
Total Costs                  
Profit using only Direct Costs                  
Profit margin % using Direct Costs only                  
Profit using Total Costs                  
Profit margin % using Total Costs                  
Answer #1
Department   Cost Pool Amount (A) Cost Driver   Total Utilization (B) Allocation Rate = A/B
Financial Services $4,000,000 Patient Revenue 62,750,000 0.0634
Facilities $7,500,000 Square feet 455,000 16.48
Housekeeping $5,600,000 Housekeeping hrs 164,000 34.15
Administration $8,500,000 Salary dollars 24,500,000 0.35
Indirect Cost Allocations Inpatient   Outpatient   Clinic   Total
Financial Services

0.0634*$18,500,000

= $1,179,283

0.0634*$32,000,000 = $2,039,841 0.0634*$12,250,000 =$780,876 $4,000,000
Facilities 16.48*210,000 = $3,460,800 16.48*180,000 = $2,966,400 16.48*65,000 = $1,071,200 $7,500,000
Housekeeping 34.15*80,000 = $2,732,000 34.15*58,000 = $1,980,700 34.15*26,000 = $887,900 $5,600,000
General Administration 0.35*6,500,000 = $2,255,102 0.35*12,500,000 = $4,336,735 0.35*5,500,000 = $1,908,163 $8,500,000
Total Indirect Costs $9,627,185 $11,323,676 $4,648,139 $25,600,000

P&L Statement:

Revenue = $62,750,000

Direct Costs = $33,250,000

Indirect costs = $25,600,000

Total costs = $58,850,000

Profits using only direct costs = Revenue-Direct costs = $62,750,000 - $33,250,000 = $29,500,000

Profit margin % using Direct Costs only = $29,500,000/$62,750,000 = 47.01%

Profit using Total Costs = Profits using only direct costs - Indirect costs = $29,500,000 - $25,600,000 = $3,900,000

Profit margin % using Total Costs = $3,900,000/$62,750,000 = 6.22%